OpenStudy (anonymous):

Solve. 2x2 + 13x = –20.

OpenStudy (anonymous):

2x2?

OpenStudy (anonymous):

x= –5 over 2 and x= –4 x = 5 and x = 4 x = –5 over 2 and x = –5 over 4 x = 1 over 3 and x = 3

OpenStudy (anonymous):

sorry i mean 2x^2

OpenStudy (anonymous):

2x^2 + 13 x + 20 = 0

OpenStudy (anonymous):

factor it

OpenStudy (anonymous):

you get (x + 4) (2x + 5) = 0 x = -4 x = -5/2

OpenStudy (anonymous):

i got x= –5/2 and x=4 is that correct?

OpenStudy (anonymous):

no it should be x = -5/2 and x = -4

OpenStudy (anonymous):

yes thank you.